OpenStudy (anonymous):

An equation which represents the average weight in pounds of an adult male of a given height in inches is w = 4.5h - 142. Predict the weight in pounds of an average male with a height of 6' 3" (75 inches). Round to the nearest pound. A. 168 B. 182 C. 196 D. 212

OpenStudy (australopithecus):

w = 4.5h - 142 h = 75 w = 4.5(75) - 142

OpenStudy (australopithecus):

you can solve for w on your on at this point just use a calculator

OpenStudy (australopithecus):

https://www.wolframalpha.com/input/?i=w+%3D+4.5%2875%29+-+142 here is the answer

OpenStudy (kropot72):

Substitute the value for height: w = (4.5 * 75) - 142 = 195.5 the nearest choice is C.